OpenStudy (anonymous):

Is there a formula for Tangents, Sines and Cosines? Or must I use a calculator/chart?

OpenStudy (tboi):

use a calculator/chart

OpenStudy (tboi):

Thank you and @soprano.h.d0816 may I please have a medal ill give you one

OpenStudy (soprano.h.d0816):

Law of Sines is: \[ \frac{ a }{ sinA } = \frac{ b }{ sinB } = \frac{ c }{ sinC }\] Law of Cosines is: \[a ^{2} = b ^{2} + c ^{2} - 2 bc cosA\] \[b ^{2} = a ^{2} + c ^{2} - 2ac cosB\] \[c ^{2} = a ^{2} + b ^{2} - 2ab cosC\]

OpenStudy (tboi):

@milliaunjiral

OpenStudy (soprano.h.d0816):

I don't know tangents...

OpenStudy (anonymous):

okay then, thanks

OpenStudy (soprano.h.d0816):

I found this on Google when I looked up "Formula for Tangent"

OpenStudy (soprano.h.d0816):

I hope that helps some(:

OpenStudy (tboi):

oh lol i guess there is a formula

OpenStudy (anonymous):

range = 13 to 20 IQR = 14.5 to 19 range = 7 IQR = 4.5 range = 20 IQR = 4 range = 7 IQR = 5

OpenStudy (tboi):

B @milliaunjiral

OpenStudy (anonymous):

ok